OpenStudy (anonymous):

Determine the prime factorization of 88. A. 2 * 3^2 * 5 B. 2^2 * 3 * 5 C. 2^3 * 11 D. 2^2 * 11

OpenStudy (amistre64):

since prime factorization is unique, which of those produce 88?

OpenStudy (unklerhaukus):

can you think you two number that multiply together to give eighty-eight?, if you can't, check your times tables

OpenStudy (anonymous):

Solve each equation A. 2* (3^2) *5 = _____ B.(2^2) *3 *5 = ____ C. (2^3) * 11 = ____ D. (2^2) * 11 = ____ Whichever gave you 88 is the answer